# Exploiting Precision Mapping and Component-Specific Feature Enhancement for Breast Cancer Segmentation and Identification

### v4

@@ 3.5.1 Convolutional Block
3.5.2 Double Convolutional Block
The double convolutional block comprises two consecutive convolutional layers with 256 filters, a kernel size of 3, and a padding of 1. Mathematically, the operation of this block can be represented as
$\underbrace{O}_{\text{Output}}=\underbrace{\text{conv}_{2}}_{\text{Second convolution}}\left(\underbrace{\text{conv}_{1}}_{\text{First convolution}}\left(\underbrace{I}_{\text{Input}}\right)\right)$ (26)
3.5.3 Self-Attention Block
The self-attention block in DCNNIMAF computes the attention weights $\alpha_{ij}$ for each pair of positions $(i,j)$ within the feature map.
$\underbrace{\alpha_{ij}}_{\text{Attention weight for position }(i,j)}=\underbrace{\text{softmax}\left(\frac{\underbrace{QK^{T}}_{\text{Dot product of query and key}}}{\sqrt{\underbrace{d_{k}}_{\text{Dimensionality of key vectors}}}}\right)}_{\text{Softmax normalization}}\underbrace{V}_{\text{Value matrix}}$ (27)
3.5.4 Workflow and Execution
The flow of information through DCNNIMAF begins with an input layer of shape (256, 256, 3). Initially, the segmentation map undergoes a convolutional block with 512 filters, a padding of 2 and a kernel size of 3. This extracts low-level features such as textures and edges from the input. Following this, the output from the first convolutional block is passed through another convolutional block with 256 filters, the same kernel size, and padding, but with SiLU activation. The introduction of SiLU activation enhances the non-linearity for higher-level feature extraction, which helps to distinguish between different breast tissue characteristics indicative of cancerous growth.
Subsequently, a double convolutional block is applied to further refine feature extraction. By employing consecutive convolutional layers with 256 filters each, this block extracts deeper and more abstract features from the input. Following this, a convolutional block with 128 filters, a kernel size of 4, and padding of 2 is employed, accompanied by a leaky ReLU activation. This operation aims to distill the extracted features into more compact and discriminative representations, facilitating the model’s capability to detect, and interpret complex patterns within the tumor’s structure such as textural anomalies to irregular shapes Continuing the feature refinement process, another convolutional block with 128 filters, a padding of 1, and a kernel size of 3 is applied, this time utilizing SiLU activation.
Subsequently, two convolutional blocks are utilized - the first with 128 filters, a kernel size of 4, and a padding of 2, and the second with 64 filters, a padding of 1, and a kernel size of 3. These features are then fed to a spatial attention mechanism, enhancing the model’s capacity to adjust to subtle differences between various tissue characteristics associated with malignant and benign tumors.
@@ 3.5.4 Workflow and Execution
The feature map obtained from the preceding operations is then concatenated with the output from a convolution and batch normalization layer with 64 filters, a padding of 2, and a kernel size of 3. This model integrates both high-level and low-level features across different layers through a concatenation approach, enabling a more comprehensive representation of the input image. This allows the model to learn about the presence of microcalcifications and the density of the tumor tissue, that are most indicative of malignancy.
This concatenated output undergoes further processing through convolutional and activation layers before being upsampled and concatenated again with intricate feature attention results. This iterative refinement process ensures that the model can effectively leverage both global and local contextual details present in the input segmentation map. This is then fed through additional convolutional blocks and pooling layers before being passed through a self-attention block. By incorporating self-attention mechanisms, it allows the model to highlight more weightage to the distribution of cells or the presence of necrosis, filtering out less relevant information and potential artifacts that could obscure diagnosis.
Ultimately, the result from the self-attention block is flattened and subjected to dropout regularization to mitigate overfitting. Dropout prevents the model from relying on specific features or patterns within the training data that may not generalize well to unseen samples, thereby improving its robustness and generalization performance.
The feature map is then directed into a fully connected layer containing 128 neurons, then proceeds to an output layer with three neurons and softmax activation for classification into malignant, benign, or normal categories. This final step consolidates the extracted features into a compact representation suitable for classification, enabling the model to make accurate predictions concerning the existence and severity of breast tumors based on the input ultrasound image. The model’s training parameters were updated after each epoch via backpropagation using the categorical cross entropy loss criterion (29).
@@
The working algorithm in classifying breast cancer as benign, malignant, or normal, is demonstrated in Algorithm 3.
The model specifications and parameters of the proposed DCNNIMAF classifier have been shown in Table 2.
Table 2: Architecture Specifications of DCNNIMAF Classification Model
@@ 4 Experimental Setup and Results
This section outlines the findings and discussion achieved from training the proposed models. The experiments were conducted in a system with the following specifications: CPU - AMD Ryzen 7 4800H with Radeon Graphics, x86_64 architecture, running at a speed of 3GHz with 8 cores; GPU - NVIDIA GeForce RTX 3050-PCI Bus 1; and 32GB of RAM. These details are summarized in Table 3.
@@ 4.1 Segmentation Evaluation Metrics
The proposed segmentation framework’s performance was evaluated during the training and validation phase using the following segmentation metrics:
@@ 4.1.4 Performance Evaluation and Discussion
From the segmentation results, it can be inferred that this model has demonstrated impressive performance. The high values obtained from IoU, Dice Coefficient, and Accuracy scores, along with the minimal total loss imply that the InceptionResNet backbone managed to successfully extract important characteristics from input preprocessed images, and the dual-attention mechanism in the decoder blocks helped fine-tune the segmentation maps during segmentation.
Grad-CAMs, which stands for Gradient-weighted Class Activation Mapping, is a method in deep learning used to visualize important regions in an input image that guide the model’s decision-making process [31]. They are particularly useful in understanding how Convolutional Neural Networks (CNNs) make their predictions, especially in tasks like medical image segmentation, where it is necessary to observe if the attention mechanism carries out its operations properly.
The GradCAMs of the attention block at the topmost decoder block, as provided in Figure 12, show how the attention mechanism focuses on specific regions of the feature map, highlighting the importance of these regions for the segmentation task. This visualization helps in understanding how the attention mechanism contributes to the segmentation performance by emphasizing the most relevant features and their spatial locations. From GradCAMs, it can be observed that the attention mechanism progressively shifts its focus towards the tumor region, with an improvement in localization accuracy as the number of training epochs increases.
@@
Proposed Spatial-Channel Attention LinkNet Framework with InceptionResNet Backbone 97.20 96.91
U-Net [32] model attains a Dice coefficient of 82.52% and an IoU score of 69.76%. These scores reflect a foundational capability in segmenting tumors from breast ultrasound images and highlight the model’s limitations in capturing the full extent of tumor boundaries and internal structures, particularly in the nuanced textures and densities often found in breast tissues. Res U-Net [33] enhances the original U-Net with a Dice coefficient of 88% and an IoU score of 80%, demonstrating enhanced performance through the incorporation of residual connections, but further refinements in its network architecture and feature extraction are necessary to achieve optimal segmentation accuracy, especially in dealing with the variable echo intensities and shadowing effects commonly encountered in breast ultrasound imaging. By integrating a DenseNet backbone, the U-Net with DenseNet Backbone [34] reaches a Dice coefficient of 89.8% and an IoU score of 79.1%, showcasing the benefits of dense connectivity in improving segmentation outcomes. However, additional strategies may be required to fully leverage the complex patterns inherent in breast ultrasound images, such as the differentiation between cystic and solid components of tumors, which is critical for accurate diagnosis. The Multi-scale Fusion U-Net [35] achieves a Dice coefficient of 95.35% and an IoU score of 91.12%, marking a significant improvement over earlier models. But it shows suboptimal performance when handling the heterogeneity of breast tissues and the dynamic nature of tumor growth observed in ultrasound sequences. The proposed Spatial-Channel Attention LinkNet Framework with InceptionResNet Backbone stands out with a Dice coefficient of 97.20% and an IoU score of 96.91%. This performance is attributed to the integration of spatial-channel attention mechanisms and the robust InceptionResNet backbone, which together enable precise localization and delineation of tumors, including the ability to distinguish between different types of breast lesions based on their texture, shape, and boundary characteristics.
@@ 4.2 Classification Evaluation Metrics
The outcomes of the proposed DCNNIMAF model for breast cancer classification are evaluated during the training and validation phase using the following classification metrics:
@@ 4.2.1 Accuracy
Accuracy is a fundamental metric that evaluates the overall performance of a model across all classes. It measures the proportion of true classifications (both true positives and true negatives) in the total images classified, providing a comprehensive view of the model’s effectiveness in correctly classifying instances.
@@
  • $TP$ denotes the number of true positives.
  • $TN$ denotes the number of true negatives.
  • $FP$ denotes the number of false positives.
  • $FN$ denotes the number of false negatives.
  • $n$ denotes the total number of classes.
@@ 4.2.3 Recall
Recall, also known as sensitivity, measures the ability of the classifier to identify all relevant instances within a specific class. It is crucial in situations where missing a positive instance (false negative) is more detrimental than identifying a negative instance as positive (false positive). Recall helps in ensuring that the model does not overlook any relevant instances.
@@ 4.2.4 F1-Score
The proposed DCNNIMAF classifier was trained for 100 epochs, and the evaluation metrics were recorded after each epoch. The training and validation curves obtained for accuracy coupled with categorical cross-entropy loss have been depicted in Figures 13 and 14 respectively. From the graph plots, it can be seen that the classification model has obtained a high accuracy of 99.2% at a minimal loss of 0.03. Figures 15, 16, and 17 display the training and validation precision, recall, and F1-score curve, respectively. It can be inferred from the graphs, that the proposed model has minimized false positives and false negatives, thereby achieving a remarkable precision of 99.3% and a recall of 99.1%. The high values of precision and recall contribute to the high F1-score value of 99.1%.
Figure 13: Training and Validation Accuracy Curves of Proposed DCNNIMAF Classifier
Figure 14: Training and Validation Loss Curves of Proposed DCNNIMAF Classifier
Figure 15: Training and Validation Precision Curves of Proposed DCNNIMAF Classifier
Figure 16: Training and Validation Recall Curves of Proposed DCNNIMAF Classifier
Figure 17: Training and Validation F1-Score Curves of Proposed DCNNIMAF Classifier
@@ 4.2.5 Performance Evaluation and Discussion
The normalized confusion matrix obtained on the validation data using the trained DCNNIMAF classification model has been presented in Figure 18. A normalized confusion matrix is a type of confusion matrix where the values are normalized to show proportions or percentages. It is useful for comparing classification performance across classes, since the values are between 0 to 1, making it easy to interpret.
Figure 18: Confusion Matrix Obtained from Proposed DCNNIMAF Classifier
@@
The proposed DCNNIMAF model is compared with other pretrained CNNs, including EfficientNetV2[36], MobileNetV2 [37], [38], NASNetMobile[39], Xception[40], InceptionV3[41], InceptionResNetV2[30], MobileNet[42], VGG16[43], and ResNet50[44]. This comparison aims to provide an overall assessment of the proposed model relative to existing baseline CNNs widely utilized for breast cancer classification. All models, including the proposed one, are trained utilizing the identical dataset, and the outcomes are presented in Table 5. The performance of these models is evaluated based on the following metrics: Accuracy (Acc), Precision (Prec), Recall (Rec), and F1 Score (F1).
@@
DCNNIMAF (Proposed) 0.989 0.994 0.992 0.993 0.992 0.993 0.991 0.991
From Table 5, it is evident that the proposed DCNNIMAF model has outperformed all baseline CNN models in terms of performance evaluation metrics. EfficientNetV2 overfits on the data due to difficulty in generalizing the nuanced features of breast cancer like irregular margins of malignant lesions or varying degrees of echogenicity observed in ultrasound images. MobileNetV2’s lightweight architecture struggles with the detailed analysis required to detect early signs of breast cancer, such as subtle changes in echotexture or the presence of microcalcifications within lesions. While DenseNet121 benefits from dense connectivity for feature reuse, its performance in identifying specific breast cancer markers like the orientation and distribution of calcifications or the assessment of lesion vascularity is compromised. NASNetMobile, designed for mobile applications, lacks the precision needed to capture the complex interplay of features indicative of breast cancer, such as the irregular shapes of masses or variations in posterior acoustic shadowing. Xception does not fully exploit the spatial dependencies crucial for identifying specific indicators of breast cancer, such as the pattern of calcifications or the echogenicity of surrounding tissue.
InceptionV3’s design compromise for computational efficiency limits its capacity to analyze the multidimensional data characteristic of breast cancer ultrasound images, particularly in detecting subtle architectural distortions or changes in tissue echotexture. Despite its sophisticated architecture, InceptionResNetV2 does not optimally align with the need to identify specific, disease-related features like the texture and margin irregularities of masses or the presence of ductal abnormalities. MobileNet’s focus on efficiency limits its depth necessary for detailed feature extraction from breast cancer ultrasound images. VGG16’s simplicity and relative shallowness struggles with the detailed analysis required to detect and classify features such as the presence of posterior acoustic enhancement, leading to lower accuracy in validation tests. Features such as the assessment of lesion margins might not be adequately learned due to limitations in the ResNet50’s depth and focus. The proposed DCNNIMAF model distinguishes itself by effectively integrating multiple spatial and self-attention mechanisms, enabling precise identification of critical features such as calcifications, architectural distortions, and mass margins. These enhancements allow the model to capture the complex, heterogeneous pathology of breast cancer evident in ultrasound imagery.
@@
From the results presented in Table 6, it is apparent that the DCNNIMAF model proposed in this research outperforms all other models in existing research. The assembly of Fine Tuned VGG16 and VGG19 [45] achieves moderate performance with accuracy and F1-scores around 95%. Its performance is relatively low, indicating potential limitations in its ability to capture the complexity of breast cancer pathology fully. CNN-based Ensemble Learner with MLP Meta Classifier [46] has shown high performance with an accuracy of 98% but has struggled with identifying subtle changes in the irregular shapes of masses. BCCNN [47] shows promising results with metrics around 98%. However, the slight variation in F1-score compared to the highest performers suggests it faces challenges in maintaining a balance between precision and recall, essential for minimizing errors in breast cancer diagnosis.
@@
ResNet50 Hybrid with SVM [48] presents strong recall but exhibits a lower precision score. This discrepancy indicates that while the model is capable of identifying many positive cases, it struggles with accurately distinguishing between benign and malignant lesions, leading to potential false positives. The precision score of Deep CNN with Fuzzy Merging [49] drops significantly highlighting a critical issue in its ability to classify breast cancer cases precisely. This suggests that while the model captures broad patterns effectively, it overlooks finer details necessary for accurate diagnosis. Xception combined with SVM R [50] shows a balanced performance of around 96% but indicates a relative inefficiency in comparison to other models in terms of feature extraction capabilities, leading to inefficiency in real-world use. Grid-based Deep Feature Generator with DNN Classifier [51] demonstrates a high precision score, but the minor discrepancies in recall and F1-score indicate potential inefficiencies in capturing all relevant pathological features, affecting its overall efficacy.
InceptionV3 with Residual Connections [52] achieves a high recall but significantly lower precision, indicating a significant imbalance in its diagnostic capabilities. This suggests challenges in accurately discriminating between similar-looking benign and malignant cases, which is crucial for reducing false positives. EDLCDS-BCDC [53] presents moderate performance across metrics, around 95% to 97%, highlighting potential shortcomings in accurately identifying subtle differences. AlexNet, ResNet50, and MobileNetV2 Hybrid Feature Extractor with mRMR and SVM [54]. shows solid performance with accuracy and F1-scores around 95%. However, its limitations suggest shortcomings in fully adapting to the complex and varied nature of breast cancer pathology, indicating areas for potential enhancement.
The proposed DCNNIMAF model demonstrates remarkable performance across all metrics evaluated, surpassing all other models in this comparison. This can be attributed to its meticulously designed architecture that incorporates advanced feature extraction techniques and multiple attention mechanisms, allowing for the precise and effective identification of the nuanced pathological features associated with breast cancer. This specialized approach ensures not only high accuracy but also maintains excellent precision and recall, showcasing its robustness and reliability in clinical applications for breast cancer classification.

### v5

@@ 3.5.1 Workflow and Execution
The feature map from the spatial attention mechanism of the CSFEM is concatenated with the output from a convolution and batch normalization (10) layer with 64 filters, padding of 2, and a kernel size of 3. This concatenation integrates both high-level and low-level features across different layers, enabling a more comprehensive representation of the input image. The enriched feature map is further processed through convolutional and activation layers before being upsampled and concatenated with intricate feature attention results. This iterative refinement ensures that the model effectively leverages both global and local contextual details present in the input segmentation map.
Following this, the feature map is flattened and subjected to dropout regularization to mitigate overfitting. Dropout prevents the model from relying on specific features or patterns within the training data that may not generalize well to unseen samples, thereby improving its robustness and generalization performance. The dropout-regularized features are then directed into a fully connected layer containing 128 neurons, which transforms the features into a compact representation suitable for classification. Finally, the output layer consists of three neurons with a softmax activation function (18) to classify the input into malignant, benign, or normal categories. The algorithmic workflow of our CSFEC-Net is provided in Algorithm 3.
Algorithm 3 CSFEC-Net for Breast Cancer Classification
0:  Segmented breast tumor images $\text{breastTumor}_{\text{segmented}}$
0:  Trained CSFEC-Net classifier model $\text{model}_{\text{classification}}$
1:
2:  Function TrainClassifier($\text{breastTumor}_{\text{segmented}}$):
3:  Initialize parameters $\theta$ of $\text{model}_{\text{classification}}$
4:  $\beta\leftarrow$ batch size
5:  $T\leftarrow$ total epochs
6:  for epoch $\leftarrow 1$ to $T$ do
7:     $\mu\leftarrow$ learning rate
8:     for each batch $(x,y)$ sampled from training data do
9:        $\hat{y}\leftarrow\text{model}_{\text{classification}}(x)$
10:        $\mathcal{L}_{\text{loss}}\leftarrow$ categorical cross-entropy loss between $y$ and $\hat{y}$
11:        Update $\theta$ using backpropagation with loss $\mathcal{L}_{\text{loss}}$:
12:        $\theta\leftarrow\theta-\mu\nabla_{\theta}\mathcal{L}_{\text{loss}}$where $\nabla_{\theta}\mathcal{L}_{\text{loss}}$ is the gradient of the loss with respect to $\theta$}
13:     end for
14:     if validation performance plateaus then
15:        Adjust learning rate $\mu$ to promote further learning
16:     end if
17:  end for
18:  $\text{output}_{\text{DCNNIMAF}}\leftarrow\text{model}_{\text{classification}}$
19:  return  $\text{output}_{\text{DCNNIMAF}}$
3.5.2 Training Setup and Specifications
Our proposed CSFEC-Net’s training parameters were updated after each epoch via backpropagation using the categorical cross entropy loss criterion (27).
@@ 3.5.2 Training Setup and Specifications
The model specifications and parameters of the proposed CSFEC-Net classifier have been shown in Table 2.
Table 2: Specifications of the Proposed CSFEC-Net Classification Model
@@ 4 Experimental Setup and Results
This section presents the results and arguments derived from the application of the proposed models. The experiment was realized on a machine with the following infrastructure: CPU - AMD Ryzen 7 4800H with Radeon Graphics, x86_64 architecture, at a speed of 3GHz with 8 cores; GPU - NVIDIA GeForce RTX 3050-PCI Bus 1; and RAM is 32 GB. These points are summarized in Table 3.
@@ 4.1 Segmentation Evaluation Metrics
The evaluations in terms of the proposed segmentation framework were carried out in the training and validation phase using the following metrics entailing segmentation:
@@ 4.1.4 Performance Evaluation and Discussion
From the segmentation results, it can be extrapolated that the said model has performed remarkably well. High values of IoU, Dice Coefficient, and accuracy scores counterbalanced by very little total losses reflect that the InceptionResNet backbone is very good at extracting valuable features out of input preprocessed images and the PMM module inserted in the decoder blocks fine-tunes the segmentation maps for segmentation [31].
Grad-CAM stands for gradient-weighted class activation mapping and is an algorithm denoting a methodology applied in DL for visualizing the different important areas that could influence a model’s decision in an input image. The primary area of this research is how the CNNs make predictions in different tasks and also applies in medical image segmentation wherein the attention mechanism must be observed if it does its operations correctly.
@@
The GradCAMs with respect to the attention block of the topmost decoder block as illustrated by Figure 12 shows us how the attention mechanism can focus on certain regions within the feature map and highlight their importance towards the segmentation task. This visualization can illustrate that the attention mechanism contributes to the segmentation performance by highlighting the most relevant features with respect to their most important spatial locations. From GradCAMs, it can be observed that the attention mechanism progressively shifts its focus towards the tumor region, with an improvement in localization accuracy as the number of training epochs increases. Table 4 compares the performances of existing literature in breast lesion segmentation against our proposed PMAD-LinkNet.
@@
PMAD-LinkNet (Proposed) 97.20 96.91
From Table 4, we can see that U-Net [32] attains a Dice coefficient of 82.52% and an IoU score of 69.76%. These scores reflect a foundational capability in segmenting tumors from breast ultrasound images and highlight the model’s limitations in capturing the full extent of tumor boundaries and internal structures, particularly in the nuanced textures and densities often found in breast tissues. Res U-Net [33] enhances the original U-Net with a Dice coefficient of 88% and an IoU score of 80%, demonstrating enhanced performance through the incorporation of residual connections, but further refinements in its network architecture and feature extraction are necessary to achieve optimal segmentation accuracy, especially in dealing with the variable echo intensities and shadowing effects commonly encountered in breast ultrasound imaging. By integrating a DenseNet backbone, the U-Net with DenseNet Backbone [34] reaches a Dice coefficient of 89.8% and an IoU score of 79.1%, showcasing the benefits of dense connectivity in improving segmentation outcomes. However, additional strategies may be required to fully leverage the complex patterns inherent in breast ultrasound images, such as the differentiation between cystic and solid components of tumors, which is critical for accurate diagnosis. The Multi-scale Fusion U-Net [35] achieves a Dice coefficient of 95.35% and an IoU score of 91.12%, marking a significant improvement over earlier models. But it shows suboptimal performance when handling the heterogeneity of breast tissues and the dynamic nature of tumor growth observed in ultrasound sequences. The proposed Spatial-Channel Attention LinkNet Framework with InceptionResNet Backbone stands out with a Dice coefficient of 97.20% and an IoU score of 96.91%. Feature analysis is carried out using spatial-channel attention mechanisms together with a robust InceptionResNet backbone, enabling precise localization and delineation of tumors, including distinguishing type of breast lesions based on texture, shape, and boundary characteristics.
@@ 4.2 Classification Evaluation Metrics
The classification performance outcomes that have been proposed by the CSFEC-Net model for breast cancer classification are measured during training and validation using the following classification metrics:
@@ 4.2.1 Accuracy
Accuracy is the basic measure for judging the performance of the model as a whole concerning all classes. It gives a proportion of true classifications (true positive and true negative) against the total number of images classified, thereby giving a picture of the overall performance of the model while classifying instances.
@@
  • $TP$ represents the number of true positives.
  • $TN$ represents the number of true negatives.
  • $FP$ represents the number of false positives.
  • $FN$ represents the number of false negatives.
  • $n$ represents the total number of classes.
@@ 4.2.3 Recall
Recall is a measure of how well the classifier can find all the relevant instances that belong in a specific class. It is important when missing a positive instance (false negative) is far worse than classifying a negative one as positive (false positive). Recall helps ensure that the model has not left any important cases out.
@@ 4.2.4 F1-Score
The proposed CSFEC-Net classifier was trained for 100 epochs, and the evaluation metrics were recorded after each epoch. The training and validation curves obtained for accuracy coupled with categorical cross-entropy loss have been depicted in Figures 14 and 14 respectively. From the graph plots, it can be seen that the classification model has obtained a high accuracy of 99.2% at a minimal loss of 0.03. Figures 16, 16, and 17 display the training and validation precision, recall, and F1-score curve, respectively. It can be inferred from the graphs, that the proposed model has minimized false positives and false negatives, thereby achieving a remarkable precision of 99.3% and a recall of 99.1%. The high values of precision and recall contribute to the high F1-score value of 99.1%.
Figure 13: Training and Validation Accuracy Curves of Proposed CSFEC-Net Classifier
Figure 14: Training and Validation Loss Curves of Proposed CSFEC-Net Classifier
Figure 15: Training and Validation Precision Curves of Proposed CSFEC-Net Classifier
Figure 16: Training and Validation Recall Curves of Proposed CSFEC-Net Classifier
Figure 17: Training and Validation F1-Score Curves of Proposed CSFEC-Net Classifier
@@ 4.2.5 Performance Evaluation and Discussion
The normalized confusion matrix obtained on the validation data using the trained CSFEC-Net classification model has been presented in Figure 18. A normalized confusion matrix is a type of confusion matrix where the values are normalized to show proportions or percentages. It is useful for comparing classification performance across classes, since the values are between 0 to 1, making it easy to interpret.
Figure 18: Confusion Matrix Obtained from Proposed CSFEC-Net Classifier
@@
The proposed CSFEC-Net model is compared with other pretrained CNNs, including EfficientNetV2[36], MobileNetV2 [37], [38], NASNetMobile[39], Xception[40], InceptionV3[41], InceptionResNetV2[30], MobileNet[42], VGG16[43], and ResNet50[44]. This comparison aims to provide an overall assessment of the proposed model relative to existing baseline CNNs widely utilized for breast cancer classification. All models, including the proposed one, are trained utilizing the identical dataset, and the results are presented in Table 5. The performance of these models is evaluated based on the following metrics: Accuracy (Acc), Precision (Prec), Recall (Rec), and F1 Score (F1).
@@
CSFEC-Net (Proposed) 0.989 0.994 0.992 0.993 0.992 0.993 0.991 0.991
From Table 5, it is evident that the proposed CSFEC-Net model has outperformed all baseline CNN models in terms of performance evaluation metrics. EfficientNetV2 overfits on the data due to difficulty in generalizing the nuanced features of breast cancer like irregular margins of malignant lesions or varying degrees of echogenicity observed in ultrasound images. MobileNetV2’s lightweight architecture struggles with the detailed analysis required to detect early signs of breast cancer, such as subtle changes in echotexture or the presence of microcalcifications within lesions. While DenseNet121 benefits from dense connectivity for feature reuse, its performance in identifying specific breast cancer markers like the orientation and distribution of calcifications or the assessment of lesion vascularity is compromised. The lack of accurate irregular mass shapes and variation in posterior acoustic shadowing are some features that characterize breast cancer that might not be picked well by NASNetMobile because it was designed simply for mobile applications. On the other hand, Xception does not take advantage of the spatial dependencies that are important in determining certain indicators of breast cancer, such as the pattern of calcifications or the echogenicity of surrounding tissue.
InceptionV3’s design compromise for computational efficiency limits its ability to analyze multidimensional data, characteristic of breast cancer ultrasound images, particularly in detecting subtle architectural distortions or changes in tissue echotexture. Despite its sophisticated architecture, InceptionResNetV2 does not optimally resonate with the need to identify very specific, disease-related features like the texture and margin irregularities of masses or the presence of ductal abnormalities. MobileNet’s efficiency does dictate how its features should not be very deep, thus failing on the level of detail required during feature extraction by breast cancer ultrasound images. The simplicity and related shallowness of VGG16 struggle with a more intricate analysis needed for detecting and classifying features such as posterior acoustic enhancement, accounting for lower accuracy in validation tests. Features such as the assessment of lesion margins might not be adequately learned due to limitations in the ResNet50’s depth and focus. The proposed model’s integrated CSFEM effectively leverages multi-level attention mechanisms for the identification of critical features such as calcifications, architectural distortions, and mass margins across diverse tumor segmentation qualities. These enhancements allow the model to capture the complex, heterogeneous pathology of breast cancer evident in ultrasound imagery.
@@
From the results presented in Table 6, it is apparent that the CSFEC-Net model proposed in this research outperforms all other models in existing research. The assembly of Fine Tuned VGG16 and VGG19 [45] achieves moderate performance with accuracy and F1-scores around 95%. Its performance is relatively low, indicating potential limitations in its ability to capture the complexity of breast cancer pathology fully. CNN-based Ensemble Learner with MLP Meta Classifier [46] has shown high performance with an accuracy of 98% but has struggled with identifying subtle changes in the irregular shapes of masses. BCCNN [47] shows promising results with metrics around 98%. However, the slight variation in F1-score compared to the highest performers suggests it faces challenges in maintaining a balance between precision and recall, essential for minimizing errors in breast cancer diagnosis.
@@
The SVM hybrid with ResNet50 [48] shows high recall but is weak in precision. Thus, while this model is good at identifying the number of positive cases, it has difficulty discerning benign lesions from malignant ones, which tends to lead it to false positives. The precision score drops markedly in the case of Deep CNN with Fuzzy Merging [49], alerting attention to a serious failure in its ability to classify breast cancer cases accurately. This implies that it might be good at identifying macro patterns but poor at capturing the important micro ones that make for accurate diagnosis. The Xception combined with SVM R [50] shows a fair performance of around 96% which points to some relative inefficiency as compared to the other models with respect to feature extraction capabilities, which makes it inefficient for realistic use. Grid-based Deep Feature Generator with DNN Classifier [51] has high precision scores, but the insignificant differences in recall and F1-score might point toward some ineffectiveness in the full capture of relevant pathological features, which will reduce its overall efficacy.
InceptionV3 with Residual Connections achieves a recall of very high values while at the same time having a precision value substantially lower than the recall, indicating a huge asymmetry in diagnostic capabilities. Such imbalances are usually due to very great discrimination difficulties in benign cases against malignant ones, which is necessary to cut down false positives. On the other hand, EDLCDS-BCDC gives a moderate report in that range of metrics, between 95% to 97%, which suggests it cannot be very accurate on spotting very minute differences. AlexNet, ResNet50, and MobileNetV2 Hybrid Feature Extractor with mRMR and SVM shows outstanding performance with an accuracy and F1-score of about 95%. This limitation of the model suggests that it is unable to fully adapt to the complex and diverse pathology of breast cancer and can be improved.
The proposed CSFEC-Net model has also outperformed all other models in this comparison with high performance in almost every metric evaluated. The key factor for this performance is its rigorous architectural design, inclusive of an exclusive CSFEM with multiscale attentions, to enable the most precise and optimal extraction of the subtle pathological features of breast cancer from the data. This individualized strategy provides both high accuracy and tremendous precision and recall and demonstrates robustness and reliability in clinical applications for breast cancer classification.